# AgileFormer: Spatially Agile Transformer UNet for Medical Image Segmentation

### v1

@@ 3.3 Ablations on model design variants
Table 4: Ablation studies on model design variants in Synapse multi-organ segmentation task, starting from a SwinUNet baseline. This is also the roadmap going from SwinUNet to the proposed AgileFormer.
@@
We conducted ablation studies on the model design variants, including patch embedding and different choices of spatially dynamic attention and positional encoding (see Table˜4). First, we showed that replacing patch merging with convolutional downsampling led to a small performance gain of $0.34\%$ in average DSC. Then, adding deformable embedding can lead to a performance gain of $1.9\%$ in DSC. Second, we ablated different choices of spatially dynamic self-attention by alternating window/ deformable/neighborhood attention. We removed positional encoding to eliminate its effect for now. Alternating window attention (WMSA) with both DMSA and NSMA led to a performance gain of $0.7\%$ (DMSA) and $0.4\%$ (NSMA), respectively. Alternating NMSA and DMSA resulted in a performance gain of $1.4\%$. Third, we brought back positional encoding, demonstrating that not all PEs can lead to a performance gain, e.g., adding APE and CPE even led to a performance drop. This is because these two PEs were not designed for irregularly sampled grids introduced by DMSA. Instead, the proposed MS-DePE improved performance by $0.8\%$. Notably, the proposed AgileFormer did not lead to a huge computational burden, with only a 1.1% increase in the number of parameters and a 15% rise in floating point operations, compared to SwinUNet.
3.4 Model scaling behavior
We compared the model scaling behavior of the proposed method with other ViT-UNets. As shown in Figure˜1(a), the proposed method demonstrated improved performance with growing model sizes for the Synapse multi-organ segmentation task, where different organs exhibit a lot of variability in shape and size. Specifically, the performance was improved by 2.15 in DSC (%) going from AgileFormer-T to AgileFormer-B, with parameters increase of approximately fourfold from 28.85M to 123.47M. We also observed that the methods with dynamic modules scaled better than those without dynamic modules, except for CoTr and SDAUT. We hypothesized this might be attributed to the fact that the main backbone of CoTr was still a CNN. Besides, deformable modules in both CoTr and SDAUT were only placed in the bottleneck.
4 Conclusion
In this paper, we proposed AgileFormer that systematically introduced spatially dynamic components (i.e., deformable patch embedding, spatially dynamic self-attention, and multi-scale deformable positional encoding) into a UNet for capturing spatially dynamic information for diverse target objects in the medical image segmentation. Extensive experiments demonstrated the effectiveness of the proposed method for a variety of medical image segmentation tasks. We aspire that the idea of systematically introducing spatially dynamic components to ViT-UNet will guide future designs on how to extract spatially dynamic representation for medical image segmentation with multiple targets.
Appendix
Figure 5: (a) CoTr is a hybrid model where Deformable DETER [37] was applied to the bottleneck. (b) SDAUT is a ViT-UNet originally proposed for medical image restoration. The encoder/decoder of SDAUT used window attention (WMSA); while DMSA was applied to the bottleneck. (c) MERIT captured dynamic feature representations using a multi-resolution and multi-stage approach, with the MaxViT [38] as the building block. (d) AgileFormer alternated NMSA and DMSA as the building block, adding novel deformable patch embedding and multi-scale deformable positional encoding.
Appendix A Comparison with other dynamic ViT-UNets
@@
CoTr. Similar to its ancestor model [37], CoTr is a hybrid model where the main backbone (encoder/decoder) is still a CNN. Multi-scale deformable attention outlined in [37] serves as a bottleneck. As a result, it can only introduce limited deformability to capture spatially dynamic feature representation, as the main feature extractor does not have spatially dynamic components. In addition, CoTr does not take full advantage of ViT to capture long-range dependencies as its main backbone is still a CNN. We hypothesize these factors may limit its scalability (see evidence in Appendix C). Instead, the proposed method uses a ViT as the main backbone (encoder/decoder), where neighborhood attention and deformable attention [15] serve as the main building blocks, which better preserves the advantage of ViT while having the additional advantage of capturing spatially localized and dynamic features.
SDAUT. Although SDAUT also uses the deformable attention mechanism [15] as the proposed method, it is primarily proposed for image restoration tasks instead of a segmentation task. The optimal architecture of SDAUT uses a combination of window attention and deformable attention, where deformable attention is applied only to the third encoder and decoder. Instead, the deformable attention is applied at every layer (encoder/decoder) of the proposed method, which facilitates the extraction of spatially dynamic features. In addition, we further introduce deformable patch embedding to fundamentally change the rigid patch embedding. We also introduce a multi-scale deformable positional encoding to handle the irregularly sampled grid introduced by the deformable attention. We found that those two novel components are as important as spatially dynamic attention to enhance the segmentation performance.
@@
MERIT. MERIT mainly takes advantage of multi-resolution information to capture spatially dynamic (mainly varying size) features, by combining the features extracted from inputs of different sizes. Instead, the proposed method handles both varying sizes as well as varying shapes by spatially dynamic attention. Whereas, we use a multi-scale deformable positional encoding to further capture multi-scale information.
Appendix B Implementation details on 3D experiments
We found that the comparison in nnFormer [20] was unfair. nnFormer and SwinUNETR take images with a size of $128\times 128\times 128$ as input; while the results reported by UNETR were achieved with inputs of a size of $96\times 96\times 96$ in our implementation. As a result, the input image size was set to $96\times 96\times 96$ for all methods. We also found that there was a huge performance gap between nnFormer and the other methods (i.e., TransUNet, TransBTS, UNETR, SwinUNETR, and CoTr), which is not consistent with the performance on other segmentation tasks. We conjectured that this performance gap might be attributed to different preprocessing, training, and evaluation procedures. Therefore, in our experiments, we retrained all the methods using the nnFormer codebase to standardize the comparison. For the proposed method, the batch size was set to 4, and the initial learning rate was set to $5\times 10^{-4}$. For the baseline methods, we followed the default hyperparameter settings in their original papers. We found out that our reimplmentation resulted in a higher performance of TransUNet, TransBTS, UNETR, SwinUNETR, and CoTr compared to results reported in nnFormer, even with a smaller size input.
[…]

### v2

@@ 4.3 Ablations on model design variants
We conducted ablation studies on the model design variants on the Synapse dataset using 2D AgileFormer, including patch embedding and different choices of spatially dynamic attention and positional encoding (see Figure 2). First, replacing patch merging with convolutional downsampling led to a small performance gain of $0.34\%$ in average DSC. Then, adding deformable embedding led to a performance gain of $1.9\%$ in DSC. Second, we ablated different choices of spatially dynamic self-attention by alternating window/deformable/neighborhood attention. We removed positional encoding to eliminate its effect for now. Alternating window attention (WMSA) with both DMSA and NSMA led to a performance gain of $0.7\%$ (DMSA) and $0.4\%$ (NSMA), respectively. Alternating NMSA and DMSA resulted in a performance gain of $1.4\%$. Third, we brought back positional encoding, demonstrating that not all PEs can lead to a performance gain, e.g., adding APE and CPE even led to a performance drop. This is because these two PEs are not designed for irregularly sampled grids introduced by DMSA. Instead, the proposed $\operatorname{MS-DePE}$ improved performance by $0.8\%$. Notably, the proposed AgileFormer did not lead to a huge computational burden, with only a 1.1% increase in the number of parameters and a 15% rise in floating point operations, compared to SwinUNet.
4.4 Model scalability
We compared the scaling behavior of the proposed method with other ViT-UNets on the Synapse dataset using 2D AgileFormer. As shown in Figure 5, the proposed method demonstrated improved performance with growing model sizes for the Synapse multi-organ segmentation task, where different organs exhibited a lot of variability in shape and size. Specifically, the performance was improved by 2.15 in DSC (%) going from AgileFormer-T to AgileFormer-B, with parameters increasing approximately fourfold from 28.85M to 123.47M. We also observed that the methods (i.e., CoTr, MERIT, and AgileFormer) can capture spatially varying representations generally scaled better than standard ViT-UNets (i.e., TransUNet and SwinUNet). However, CoTr did not scale as well as MERIT and AgileFormer. We hypothesized this might be attributed to the fact that the main backbone of CoTr was still a CNN. The deformable module in CoTr was only placed in the bottleneck.
Figure 6: Visual comparison of (a) multi-organ, (b) cardiac, and (c) brain tumor segmentation across Synapse, ACDC, and Decathlon datasets. AgileFormer showed the best qualitative segmentation quality across all three segmentation tasks.
4.5 Discussion
Failure cases: Although AgileFormer demonstrated improvements over previous state-fo-the-art segmentation methods quantitatively and qualitatively, it still has limitations on highly irregular and unstructured objects (e.g., on Pancreas and Gallbladder (see Figure S3) and some tissue classes in brain tumor segmentation (see Figure S4)). This suggests that there is still room for improvement in future work with more powerful spatially varying components; whereas our investigation revealed the importance of incorporating spatially varying components into UNet designs.
Potential trade-offs between different organs: First, we observed that AgileFormer showed more promising results on challenging small and irregular organs (e.g., Gallbladder Aorta, and Pancreas) but also achieved the best performance on most large organs (see Table 1). Second, we observed that many ViT-UNets can perform equally well on the larger organs (e.g., Liver and Spleen). However, the pure CNN-based methods did not work well on the larger organs. We conjectured that this phenomena was attributed to the fact that the self-attention mechanism in ViT-UNets better capture global context than convolution in CNNs (which is typically limited by the receptive field). Third, CNN-based models (U-Net [30] and Att-UNet [26]) biased toward small organs (e.g., Aorta) due to the inherent locality of convolution in capture small objects. In contrast, the rigid patching embedding in ViT-UNets may fail to capture such localized structures and hence showed inferior performance on small organs. Notably, our investigation revealed that replacing the rigid patch embedding with a deformable patch embedding dramatically improved the performance of ViT-UNets on segmenting smaller organs (e.g., Aorta and Pancreas). Furthermore, our novel multi-scale deformable positional encoding improved the DSC (%) by 4.21% for Gallbladder segmentation.
5 Conclusion
In this paper, we proposed AgileFormer that introduced spatially dynamic components (i.e., deformable patch embedding, spatially dynamic self-attention, and multi-scale deformable positional encoding) into a standard ViT-UNet for capturing spatially dynamic information for diverse target objects in the medical image segmentation. Extensive experiments demonstrated the effectiveness of the proposed method for a variety of medical image segmentation tasks. In addition, our AgileFormer significantly outperformed recent 2D and 3D segmentation methods, setting a new state-of-the-art performance. We aspire that the idea of systematically introducing spatially dynamic components to ViT-UNet will guide future designs on how to extract spatially dynamic representations for medical image segmentation with multiple targets, which involve heterogeneous appearances.
Appendix
Table S1: Optimal training hyperparameter and model configurations for different experimental settings.
training hyperparameters model configuratoins
Models learning rate epochs batch size weight balance parameter $\lambda$ first deformable patch embedding strides downsampling patch embedding strides input size
2D AgileFormer-B [Synapse multi-organ] 2e-4 400 24 0.6 [2, 2], [2, 2] [2, 2], [2, 2], [2, 2] $224\times 224$
AgileFormer-B [ACDC cardiac] 4e-4 400 24 0.6 [2, 2], [2, 2] [2, 2], [2, 2], [2, 2] $224\times 224$
3D AgileFormer-T [Synapse multi-organ] 2e-4 1000 4 0.5 [1, 2, 2], [2, 2, 2] [2, 2, 2], [2, 2, 2], [2, 2, 2] $64\times 128\times 128$
AgileFormer-T [ACDC cardiac] 3e-4 1000 4 0.5 [1, 2, 2], [2, 2, 2] [1, 2, 2], [1, 2, 2], [2, 2, 2] $14\times 160\times 160$
AgileFormer-T [Decathlon brain tumor] 5e-4 200 4 0.5 [2, 2, 2], [2, 2, 2] [2, 2, 2], [2, 2, 2], [2, 2, 2] $96\times 96\times 96$
Appendix A Comparison to CoTr [41] and MERIT [29]
Figure S1: (a) CoTr is a hybrid model where Deformable DETER [49] is only applied to the bottleneck. (b) MERIT captures dynamic feature representations using a multi-resolution and multi-stage approach, with the MaxViT [35] as the building block. (c) AgileFormer alternates NMSA and DMSA as the building block, adding novel deformable patch embedding and multi-scale deformable positional encoding.
@@ Appendix A Comparison to CoTr [41] and MERIT [29]
Table S2: Ablation studies on model design variants in Synapse multi-organ segmentation task, starting from a SwinUNet baseline. This is also the roadmap going from SwinUNet to the proposed AgileFormer.
@@
A.1 Comparison to CoTr [41]
@@
Similar to its ancestor model [49], CoTr is a hybrid model where the main backbone (encoder/decoder) is still a CNN. Multi-scale deformable attention outlined in [49] serves as a bottleneck (see Figure S1 (a)). As a result, it can only introduce limited deformability to capture spatially dynamic feature representations, as the main feature extractor does not have spatially dynamic components. In addition, CoTr does not take full advantage of ViT to capture long-range dependencies as its main backbone is still a CNN. We hypothesized that these factors may limit its performance and scalability. Instead, the proposed method uses a pure ViT as the main backbone (encoder/decoder), where neighborhood attention and deformable attention [39] serve as the main building blocks (see Figure S1 (c)). This better preserves the advantage of ViT while having the additional advantage of capturing spatially localized and dynamic features. This accounts for the superior performance of AgileFormer compared to CoTr. In particular, AgileFormer outperforms CoTr by 5.03, 1.24, and 1.6 in DSC (%) for multi-organ, cardiac, and brain tumor segmentations, respectively.
A.2 Comparison to MERIT [29]
@@
MERIT mainly takes advantage of multi-resolution inputs to capture representation at multiple scales (see Figure S1 (b)). This is achieved by combining the features extracted from inputs of different sizes. As a result, MERIT mainly models targets with varying sizes but may fail to capture objects with varying shapes. Instead, our AgileFormer handles both varying sizes as well as varying shapes by deformable patch embedding and spatially dynamic attention (see Figure S1 (c)). In addition, we proposed a multi-scale deformable positional encoding to further capture multi-scale information. Consequently, AgileFormer outperforms MERIT by 0.84 and 0.23 in DSC (%) for multi-organ and cardiac segmentation, respectively.
More comparisons: We would like to point out that the official implementation  [^2] https://github.com/SLDGroup/MERIT of MERIT used test set to select the best model on Synapse. In contrast, we did not use the test set for model selection and only evaluated the model checkpoint at the last training epoch. As stated in Table 1, if using the test set to select the best model, Agileformer can achieve an average DSC (%) of 86.11 (vs. 84.90 MERIT). To be more rigorous, if evaluated on the last training checkpoint, MERIT only achieved a DSC (%) of 82.89 (vs. 85.74 AgileFormer) on Synapse. Therefore, the improvement over MERIT is not marginal but statistically significant.
Appendix B Additional Implementation Details
The model configurations and optimal training hyperparameter settings for 2D and 3D experiments are shown in Table S1. We used different downsampling strides for patch embeddings to accommodate different input sizes. The weight parameters $\lambda$ balance the DSC loss ($\mathcal{L}_{\text{DSC}}$) and cross-entropy loss ($\mathcal{L}_{\text{cross-entropy}}$) as follows:
$\mathcal{L}=\lambda\mathcal{L}_{\text{DSC}}+(1-\lambda)\mathcal{L}_{\text{cross-entropy}}.$ (4)
Appendix C Ablation studies
Here, we provide detailed results for ablation studies in Table S2, including the change of segmentation accuracy, parameters, and floating points (FLOPs).
Appendix D Additional Qualitative Results
Additional qualitative results on Synapse produced by 3D models are shown in Figure S2, where AgileFormer showcased accurate segmentation compared to other state-of-the-art methods. Some failure cases on Synapse are shown in Figure S3, where even methods that can extract spatially dynamic features (including AgileFormer) cannot accurately segment highly irregular organs such as Gallbladder and Pancreas. Similarly, Figure S4 shows failure cases on brain tumor segmentation tasks, where tissue classes with highly irregular appearance cannot be accurately segmented by most methods, including AgileFormer.
Figure S2: Qualitative comparison between the proposed AgileFormer, nnFormer [47], nnUNet [17], and UNETR [11] on the Synapse multi-organ segmentation task.
Figure S3: Failure cases where accurately segmenting irregular organs like the gallbladder and pancreas (highlighted by gray boxes) is challenging for most methods with spatially dynamic components [15, 29, 41], including AgileFormer.
Figure S4: Failure cases where accurately segmenting irregular tumor tissues (highlighted by gray boxes) is challenging for most methods [17, 32, 5, 41, 11, 33, 10, 47], including AgileFormer.
Appendix E Statistical Test
On Synapse, we performed the Wilcoxon signed rank test to the second-best model (i.e., 2D MERIT and 3D nnUNet), where the populations are the DSC values for 8 different organs. However, on ACDC and Decathlon datasets, we performed the paired t-test to the second-best model, as there are only three foreground labels. In these two cases, the populations were the DSC values for all subjects across the dataset.